# All-Electron Molecular Tunnel Ionization Based on the Weak-Field Asymptotic Theory in the Integral Representation

### v1

@@ Abstract
Tunnel ionization (TI) underlies many important ultrafast processes, such as high-harmonic generation and strong-field ionization. Among the existing theories for TI, many-electron weak-field asymptotic theory (ME-WFAT) is by design capable of accurately treating many-electron effects in TI. An earlier version of ME-WFAT relied on an accurate representation of the asymptotic tail of the orbitals, which hindered its implementation in Gaussian-basis-set-based quantum chemistry programs. In this work, we reformulate ME-WFAT in the integral representation, which makes the quality of the asymptotic tail much less critical, hence greatly facilitating its implementation in standard quantum chemistry packages. The integral reformulation introduced here is therefore much more robust when applied to molecules with arbitrary geometry. We present several case studies, among which is the CO molecule where some earlier theories disagree with experiments. Here, we find that ME-WFAT produces the largest ionization probability when the field points from C to O, as experiments suggest. An attractive feature of ME-WFAT is that it can be used with various types of multi-electron methods whether of density functional [Phys. Rev. A 106, 052211 (2022)] or multiconfiguration types, which facilitates tunnel ionization calculation in systems exhibiting a strong multireference character, such as transition-metal-containing molecules.
@@ I Introduction
There have been a number of efforts to develop approximate quantum mechanical theories of TI over the years, including the Ammosov-Delone-Krainov (ADK) model [7], the extension of ADK for molecular cases (MOADK) [8], the Keldysh-Faisal-Reiss (KFR) method, and the weak-field asymptotic theory (WFAT) [9, 10, 11, 12, 13] Alternatively, in direct numerical simulation TI can be evaluated using complex absorbing boundary in time-dependent wave function propagation methods [14, 15, 16, 17, 18, 19]. Among these theoretical methods, time-dependent evolution of either wave functions or densities [20, 21] are considered to be the most accurate. These methods, however, are often very expensive since at each time-step, one needs to construct the Hamiltonian from the two-electron integrals, which may reach more than several million for even medium sized molecules and moderate-sized basis sets. The alternatives to time-dependent methods are much cheaper at the expense of somewhat reduced accuracy. Of particular interest here, is the WFAT, which has been shown to give satisfactory agreement with experimental orientation-dependent ionization yields [22].
@@
Several flavors of WFAT have been developed, as detailed in the next section. Briefly, the one-electron (OE-)WFAT [9, 10, 12, 23] approximates the ionizing wave function with a single configuration framework while multi-configuration effects can be treated in the many-electron (ME-)WFAT [24, 25, 26]. In WFAT, the ionization rate can be calculated using either the asymptotic portion of the wave function, called the tail representation (TR) [9, 10, 24, 25, 26], or reformulated in the so-called integral representation (IR) [23, 27, 28]. In this work, we develop a formulation of ME-WFAT that works in virtually any molecular geometry by reformulating the theory in the integral representation (IR). Specifically, by reformulating the Schrödinger equation in its integral form via the appropriate Green’s function, the resulting integral formulas allow Gaussian-type orbitals (GTO) to be used to expand the molecular orbitals without a significant loss of accuracy. This greatly reduces the requirement on the accuracy of the asymptotic tail, and is the reason for the more general applicability of the IR formulation. Thus far, the IR has only been applied in the OE-WFAT formulation [23], where angle-dependent TI from molecules as big as benzene, naphtalene [28], and phenyl rings system [29] have been successfully calculated. IR ME-WFAT therefore paves the way towards an efficient computational method of TI that is applicable regardless of molecular geometry and is also capable of treating the effects due to multi-electron interactions. The IR ME-WFAT method presented in this work is the foundation of the DFT-based ME-WFAT in Ref. [30], where we showed that the resulting angle-dependent ionization yields are in excellent agreement with the full real-time time-dependent density functional theory calculations.
@@
We illustrate IR-MEWFAT in a range of atoms, and small- to mid-size molecular systems. We show that the main way ME-WFAT improves angle-dependent ionization rates calculated by OE-WFAT is through the definition of the dipole moment used in the formulation and the use of Dyson orbital as the ionizing orbital. In particular, we show that angle-dependent ionization rates that are in good agreement with results from experiments or time-dependent methods in CO and N2 can be recovered by using ME-WFAT. To demonstrate the versatility of the integral formulation for arbitrary molecular geometry, we demonstrate the computation of angle-dependent ionization rates in several polyatomic molecules, such as formic acid and CH3F. The computer codes used for running all simulations in this work, both ME-WFAT and OE-WFAT, are currently implemented in a developer version of the NWChem quantum chemistry program [31].
@@ II WFAT background
The earliest version of WFAT is the OE-WFAT, which is based on the single-active electron approximation [9, 10, 12]. Here, the defining problem starts from a one-electron, mean-field-type Schrödinger equation. If using Hartree-Fock (HF), this is the Roothan equation, while if density functional theory (DFT) is used, this is the Kohn-Sham equation. Choosing the defining eigenproblem this way has two consequences: (i) Only one electron is involved in the rearrangement following the ionization that forms the final state (the orbitals other than the ionizing one are frozen during ionization), and (ii) OE-WFAT can only be used in conjunction with single-determinant methods. This is because the orbital energy, needed in the calculation of ionization probability, is only well-defined for single-configuration wave functions. As a consequence of this, OE-WFAT can not simulate, even approximately, ionization starting from an excited initial state. To see why, recall that in a single-configuration framework, an excited state would be a configuration where an electron is moved from an occupied orbital to a virtual one. If, for example, the final state is the ground state (in which all electrons reside in the occupied orbitals), then the ionizing orbital is a virtual orbital of the initial charge state. Since virtual orbitals often have positive orbital energy, this violates the underlying assumption in WFAT that the initial energy should be negative.
@@ III Theory
III.1 The first asymptotic expansion of multi-electron wave functions
@@
Since we are considering single ionization, it is useful to use an ansatz in which the eigenfunctions of Eq. (1) are expanded in terms of a basis that factorizes into an antizymmetrized product between an $(N-1)$-electron function and a one-electron function,
@@
III.2 The second asymptotic expansion of multi-electron wave functions
@@
We note that the $O(r^{-2})$ terms in Eq. (18) and (19) are identical, and are incorporated to account for the dipole term of $V^{(N)}(\mathbf{R}_{N})$. For the same argument regarding the the $O(r^{-2})$ term in Eq. (13), this higher multipole term will be omitted in this work. One might notice that the various potential terms in Eq. (19)-(21) are not entirely origin-independent. This is because of the multipole expansion of the core potential, which gives rise to the origin-dependent terms, such as the $Z_{c}/r$ and the dipole potential term contained in $O(r^{-2})$. As has been shown in Ref. [10, 27], for wave functions having the exact asymptotic behavior, the structure factor and hence the ionization rate are origin-independent (see also Section III.4).
@@ III.3 The leading order weak-field approximation
Despite having obtained a closed form expression for $f_{\nu\sigma}^{n^{\prime}n}$, the various terms inside the curly brackets in Eq. (III.2), such as the initial and final wave functions, among others, are still not yet straightforwardly calculable because these terms are associated with a Hamiltonian that has an external DC field term. In addition, the corresponding Schrödinger equation for general multi-electron systems cannot be easily solved due to the need of the continuum eigensubspace of the unperturbed Hamiltonian. This is the point where the weak-field approximation is invoked.
@@
where $\beta_{\nu}^{n^{\prime}n(0)}=Z_{c}-\varkappa\left(n_{\xi}+\frac{|m|+1}{2}\right)$ and $\mu_{z}^{n^{\prime}n}$ is the lab-frame $z$-component of the vector $\bm{\mu}_{n}^{(N)}-\bm{\mu}_{n^{\prime}}^{(N-1)}$. This appearance of the dipole moment difference in ME-WFAT within the integral representation can also be found in the tail representation [24]. We note that the factor $(4\varkappa^{3}/F)^{\beta_{\nu}^{n^{\prime}n(0)}/\varkappa}$ in Eq. (39) is different from that in Eq. (50) in Ref. [23] in which the power of $\varkappa$ is $2$, instead of $3$. As can be checked by rederiving Eq. (59) in the aforementioned paper, the correct power of $\varkappa$ is $3$ (as appears in Eq. (39) above).
@@ III.4 The asymptotic cancellation problem
In IR ME-WFAT, the equation used to calculate the asymptotic coefficient (Eq. (III.3) or (43)) involves integration over the entire $4N$ space-spin coordinates, in contrast to the TR ME-WFAT version in Eq. (III.4) that leaves $\eta$ unintegrated. The fact that these integrals in IR ME-WFAT are over all space means that the asymptotic part of the Dyson orbital is no longer crucial in determining the convergence of the integration, which further implies that Gaussian standard basis functions can be used to span the orbital space making up the multi-electron wave functions. It is also for the same reason that, when the one-electron approximation is sufficient, the IR OE-WFAT [23, 27, 28] is more practical than the TR version for polyatomic molecular tunnel ionization problems. Lastly, although it is still a topic for future studies, the use of CI wave functions will be much more manageable in IR ME-WFAT compared to TR ME-WFAT [26].
@@ III.6 Application to single-determinant wave functions
It is important to note that no assumptions are imposed on the unperturbed wave functions $|\Psi_{n}\rangle$ and $\big{|}\Psi_{n^{\prime}}^{+}\big{\rangle}$ appearing in Eq. (III.3) other than being an eigenstate of the respective all-electron Hamiltonian—both can take any structure, whether single (Hartree-Fock or DFT [30]) or multi-determinant. Since determinants are the building blocks of many correlated wave function methods, in this section, IR ME-WFAT will be applied to the calculation of the asymptotic coefficient when both the initial and final wave functions are a single (Slater) determinant.
@@
Then, the integrals over the first $N-1$ electron coordinates as well as $s_{N}$ are carried out analytically while the one over $\mathbf{r}_{N}$ is done numerically using numerical quadrature. The former involves extensive determinant algebra and thus we will only cite the results of Appendix A here. By first defining $\Omega_{lm^{\prime}}^{\nu}(\mathbf{r})=\tilde{R}_{\nu l}(r)Y_{lm^{\prime}}(\theta,\varphi)$, the two-electron component reads
@@ IV Results and Discussion
where $i_{\textrm{max}}$ and $i_{\textrm{2nd max}}$ are the indices of the largest and second largest expansion coefficients of the Dyson orbital, respectively (see Eq. (63) and (59)). This ratio may be interpreted as the degree of similarity of the Dyson orbital to one of the occupied molecular orbitals of the initial state. The other component of our analysis is the dipole moment, because the dipole moments used in OE-WFAT and in ME-WFAT may have different signs and as will be shown below, this affects the relative magnitude of the ionization rates between the parallel and anti-parallel directions. But before presenting the results for molecules, we will first discuss the equivalence between TR and IR ME-WFAT in Section IV.1 using several atoms. The neutral and cation wave functions employed in all of the results presented here are obtained from a HF calculation, and the formulas derived in Section III.6 have been used.
@@ IV.1 Atoms
In Table 2, Li and Na present an exception in which their $g_{00}$ and $g_{\textrm{ref}}$ are essentially identical although HF wave functions are used. The reason is that, as can be seen from their values of $R$, which is unity, the HF procedure for the cation of these atoms produces orbitals that are almost identical to those making up the neutral HF wave function, that is, the cation is almost unrelaxed. In that case, both TR and IR ME-WFAT reduces to OE-WFAT.
@@
Fig. 1(a) shows the structure factor for the parabolic channel $(n,m)=(0,0)$ and $(n,m)=(0,1)$ of O2 calculated using OE-WFAT and a bond length of $1.158$. The orbital used for this calculation is the HOMO, which is doubly degenerate and has $\pi_{g}$ symmetry. The symmetry is reflected in the shape of the structure factor $|G_{00}|^{2}$ exhibiting two peaks at $\beta=42^{\circ}$ and $\beta=138^{\circ}$, and symmetry around $\beta=90^{\circ}$. The ionization yield of O2 interacting with an oscillating electric field of low intensity simulated using TDCI in Ref. [40] has a very similar double-peak structure located at $45^{\circ}$ and $135^{\circ}$. The shape as well as the position of these peaks also agrees with the experiments on O2 reported in Ref. [41]. The ME-WFAT results shown in Fig. 1(b) are very similar to the one-electron counterpart in panel (a) for both parabolic channels shown, with the only difference being the magnitude. In fact, the $r$ ratio for O2 obtained by HF is zero, with the largest coefficient belonging to HOMO, making the O2 Dyson orbital essentially the HOMO itself. The basis set, the ratio $r$, the largest and second largest contributors to the Dyson orbital, and the ionization potentials used in ME-WFAT or OE-WFAT for O2 as well as other molecules studied here are summarized in Table 3
Table 3: The basis set, the $r$ ratio (Eq. (53)), the largest and 2nd largest neutral’s molecular orbital, and ionization potential (IP) of the molecules studied in this work. The IP’s are in hartree, and the values to the left and right of the slash sign is the IP for ME-WFAT and OE-WFAT, respectively.

### v2

@@ Abstract
Tunnel ionization (TI) underlies many important ultrafast processes, such as high-harmonic generation and strong-field ionization. Among the existing theories for TI, many-electron weak-field asymptotic theory (ME-WFAT) is by design capable of accurately treating many-electron effects in TI. An earlier version of ME-WFAT relied on an accurate representation of the asymptotic tail of the orbitals, which hindered its implementation in Gaussian-basis-set-based quantum chemistry programs. In this work, we reformulate ME-WFAT in the integral representation, which makes the quality of the asymptotic tail much less critical, hence greatly facilitating its implementation in standard quantum chemistry packages. The integral reformulation introduced here is therefore much more robust when applied to molecules with arbitrary geometry. We present several case studies, among which is the CO molecule where some earlier theories disagree with experiments. Here, we find that ME-WFAT produces the largest ionization probability when the field points from C to O, as experiments suggest. An attractive feature of ME-WFAT is that it can be used with various types of multielectron methods whether of density functional [Phys. Rev. A 106, 052211 (2022)] or multiconfiguration types, this in turn facilitates tunnel ionization calculation in systems exhibiting a strong multireference character.
@@ I Introduction
There have been a number of efforts to develop approximate quantum mechanical theories of TI over the years, including the Ammosov-Delone-Krainov (ADK) model [7], the extension of ADK for molecular cases (MOADK) [8, 9], the Keldysh-Faisal-Reiss (KFR) method [2, 10, 11, 12], and the weak-field asymptotic theory (WFAT) [13, 14, 15, 16, 17] Alternatively, in time-dependent wave function propagation methods [18, 19, 20, 21, 22, 23], TI can be evaluated using absorbing masks [6], complex absorbing potentials [24, 25, 26, 27, 28, 29, 30], exterior complex scaling (ECS) [31, 32], infinite-range ECS [33, 34], or by projecting to scattering states [35, 36, 37, 38, 39] . Among these theoretical methods, time-dependent evolution of either wave functions or densities [40, 41] are considered to be the most accurate. These methods, however, are often very expensive since at each time-step, one needs to construct the Hamiltonian from the two-electron integrals, which may reach more than several million for even medium-sized molecules and moderate-sized basis sets. The alternatives to time-dependent methods are much cheaper at the expense of somewhat reduced accuracy. Of particular interest here, is the WFAT, which has been shown to give satisfactory agreement with experimental orientation-dependent ionization yields [42, 43, 44, 45].
@@
Several flavors of WFAT have been developed, as detailed in the next section. Briefly, the one-electron (OE-)WFAT [13, 14, 16, 46] approximates the ionizing wave function with a single configuration framework while multiconfiguration effects can be treated in the many-electron (ME-)WFAT [47, 48, 49]. In WFAT, the ionization rate can be calculated using either the asymptotic portion of the wave function, called the tail representation (TR) [13, 14, 47, 48, 49], or reformulated in the so-called integral representation (IR) [46, 50, 51]. In this work, we develop a formulation of ME-WFAT that works in virtually any molecular geometry by reformulating the theory in the integral representation (IR). Specifically, by reformulating the Schrödinger equation in its integral form via the appropriate Green’s function, the resulting integral formulas allow Gaussian-type orbitals (GTO) to be used to expand the molecular orbitals (MO) without a significant loss of accuracy. This greatly reduces the requirement on the accuracy of the asymptotic tail, and is the reason for the more general applicability of the IR formulation. Thus far, the IR has only been applied in the OE-WFAT formulation [46], where angle-dependent TI from molecules as big as benzene, naphtalene [51], and phenyl rings system [52] have been successfully calculated. IR ME-WFAT therefore paves the way towards an efficient computational method of TI that is applicable regardless of molecular geometry and is also capable of treating the effects due to multielectron interactions. The IR ME-WFAT method presented in this work is the foundation of the DFT-based ME-WFAT in Ref. [53], where we showed that the resulting angle-dependent ionization yields are in excellent agreement with the full real-time time-dependent density functional theory calculations.
@@
We illustrate IR-MEWFAT in a range of atoms, and small- to mid-size molecular systems. We show that the main way ME-WFAT improves angle-dependent ionization rates calculated by OE-WFAT is through the definition of the dipole moment used in the formulation and the use of Dyson orbital as the ionizing orbital. In particular, we show that angle-dependent ionization rates that are in good agreement with results from experiments or time-dependent methods in CO and N2 can be recovered by using ME-WFAT. To demonstrate the versatility of the integral formulation for arbitrary molecular geometry, we demonstrate the computation of angle-dependent ionization rates in several polyatomic molecules, such as formic acid and CH3F. The computer codes used for running all simulations in this work, both ME-WFAT and OE-WFAT, are currently implemented in a developer version of the NWChem quantum chemistry program [54, 55]. Another program that provides OE-WFAT capabilities is PyStructureFactor [56], while the SLIMP package is developed to target more general properties of strong-field interactions with its ionization functionality implemented at the MOADK and KFR levels of theory [57, 58].
@@ II WFAT background
The earliest version of WFAT is the OE-WFAT, which is based on the single-active electron approximation [13, 14, 16]. Here, the defining problem starts from a one-electron, mean-field-type Schrödinger equation, for example, the Roothan equation in Hartree-Fock (HF) or the Kohn-Sham equation in density functional theory (DFT). Choosing the defining eigenproblem this way has two consequences: (i) Only one electron is involved in the rearrangement following the ionization that forms the final state (the orbitals other than the ionizing one are frozen during ionization), and (ii) OE-WFAT can only be used in conjunction with single-determinant methods. This is because the orbital energy, needed in the calculation of ionization probability, is only well-defined for single-configuration wave functions. As a consequence of this, OE-WFAT can not simulate, even approximately, ionization starting from an excited initial state. To see why, recall that in a single-configuration framework, an excited state would be a configuration where an electron is moved from an occupied orbital to a virtual one. If, for example, the final state is the ground state (in which all electrons reside in the occupied orbitals), then the ionizing orbital is a virtual orbital of the initial charge state. Since virtual orbitals often have positive orbital energy, this violates the underlying assumption in WFAT that the initial energy should be negative.
@@ III Theory
III.1 The first asymptotic expansion of multielectron wave functions
@@
Since we are considering single ionization, it is useful to use an ansatz in which the eigenfunctions of Eq. (1) are expanded in terms of a basis that factorizes into an antisymmetrized product between an $(N-1)$-electron function and a one-electron function,
@@
III.2 The second asymptotic expansion of multielectron wave functions
@@
We note that the $O(r^{-2})$ terms in Eq. (18) and (19) are identical, and are incorporated to account for the dipole term of $V^{(N)}(\mathbf{R}_{N})$. For the same argument regarding the $O(r^{-2})$ term in Eq. (13), this higher multipole term will be omitted in this work. One might notice that the various potential terms in Eq. (19)-(21) are not entirely origin-independent. This is because of the multipole expansion of the core potential, which gives rise to the origin-dependent terms, such as the $Z_{c}/r$ and the dipole potential term contained in $O(r^{-2})$. As has been shown in Ref. [14, 50], for wave functions having the exact asymptotic behavior, the structure factor and hence the ionization rate are origin-independent (see also Section III.4).
@@ III.3 The leading order weak-field approximation
Despite having obtained a closed form expression for $f_{\nu\sigma}^{n^{\prime}n}$, the various terms inside the curly brackets in Eq. (III.2), such as the initial and final wave functions, among others, are still not yet straightforwardly calculable because these terms are associated with a Hamiltonian that has an external DC field term. In addition, the corresponding Schrödinger equation for general multielectron systems cannot be easily solved due to the need of the continuum eigensubspace of the unperturbed Hamiltonian. This is the point where the weak-field approximation is invoked.
@@
where $\beta_{\nu}^{n^{\prime}n(0)}=Z_{c}-\varkappa\left(n_{\xi}+\frac{|m|+1}{2}\right)$ and $\mu_{z}^{n^{\prime}n}$ is the lab-frame $z$-component of the vector difference $\bm{\mu}_{n}^{(N)}-\bm{\mu}_{n^{\prime}}^{(N-1)}$. This appearance of dipole moment difference in ME-WFAT within the integral representation is consistent with the tail representation [47], and can also be found in other TI theories that incorporate Stark shift correction [64, 65, 66]. We note that the factor $(4\varkappa^{3}/F)^{\beta_{\nu}^{n^{\prime}n(0)}/\varkappa}$ in Eq. (39) is different from that in Eq. (50) in Ref. [46] in which the power of $\varkappa$ is $2$, instead of $3$. As can be checked by rederiving Eq. (59) in the aforementioned paper, the correct power of $\varkappa$ is $3$ (as appears in Eq. (39) above).
@@ III.4 The asymptotic cancellation problem
In IR ME-WFAT, the equation used to calculate the asymptotic coefficient (Eq. (III.3) or (43)) involves integration over the entire $4N$ space-spin coordinates, in contrast to the TR ME-WFAT version in Eq. (III.4) that leaves $\eta$ unintegrated. The fact that these integrals in IR ME-WFAT are over all space means that the asymptotic part of the Dyson orbital is no longer crucial in determining the convergence of the integration, which further implies that Gaussian standard basis functions can be used to span the orbital space making up the multielectron wave functions. It is also for the same reason that, when the one-electron approximation is sufficient, the IR OE-WFAT [46, 50, 51] is more practical than the TR version for polyatomic molecular tunnel ionization problems. Lastly, although it is still a topic for future studies, the use of CI wave functions will be much more manageable in IR ME-WFAT compared to TR ME-WFAT [49].
@@ III.6 Application to single-determinant wave functions
It is important to note that no assumptions are imposed on the unperturbed wave functions $|\Psi_{n}\rangle$ and $\big{|}\Psi_{n^{\prime}}^{+}\big{\rangle}$ appearing in Eq. (III.3) other than being an eigenstate of the respective all-electron Hamiltonian—both can take any structure, whether single (Hartree-Fock or DFT [53]) or multideterminant. Since determinants are the building blocks of many correlated wave function methods, in this section, IR ME-WFAT will be applied to the calculation of the asymptotic coefficient when both the initial and final wave functions are a single (Slater) determinant.
@@
Then, the integrals over the first $N-1$ electron coordinates as well as $s_{N}$ are carried out analytically while the one over $\mathbf{r}_{N}$ is done numerically using numerical quadrature. The former involves extensive determinant algebra, and thus we will only cite the results of Appendix A here. By first defining $\Omega_{lm^{\prime}}^{\nu}(\mathbf{r})=\tilde{R}_{\nu l}(r)Y_{lm^{\prime}}(\theta,\varphi)$, the two-electron component reads
@@ IV Results and Discussion
where $i_{\textrm{max}}$ and $i_{\textrm{2nd max}}$ are the indices of the largest and second-largest expansion coefficients of the Dyson orbital, respectively (see Eq. (63) and (59)). This ratio may be interpreted as the degree of similarity of the Dyson orbital to one of the occupied MO of the initial state. For general multiconfiguration wave functions, two factors influence $r$: (1) relaxation effect, and (2) multireference effect. In the present work where all states are single-determinant, only (1) applies. The other component of our analysis is the dipole moment because the dipole moments used in OE-WFAT and in ME-WFAT may have different directions and as will be shown below, this affects the relative magnitude of the ionization rates between the parallel and antiparallel directions.
Some of the systems considered below are doubly degenerate, these are O2, OCS, and CH3F. In the exact treatment, their degeneracies will be broken by the ionizing field, however, in the leading order approximation of WFAT assumed in this work, only field-free wave functions and orbitals are used with any degeneracies still intact. It is only when one goes beyond the leading order approximation, should the field-induced modification on the orbitals be taken into account, possibly removing any degeneracies [15, 62]. The neutral and cation wave functions employed in all of the results are obtained from a HF calculation, thus representing a ground-state-to-ground-state ionization channel, and the formulas derived in Section III.6 have been used.
@@ IV.1 Atoms
In Table 2, Li and Na present an exception in which their $g_{00}$ and $g_{\textrm{ref}}$ are essentially identical, although HF wave functions are used. The reason is that, as can be seen from their values of $R$, which is unity, the HF procedure for the cation of these atoms produces orbitals that are almost identical to those making up the neutral HF wave function, that is, the cation is almost unrelaxed. In that case, both TR and IR ME-WFAT reduces to OE-WFAT.
@@
Fig. 1(a) shows the structure factor for the parabolic channel $(n,m)=(0,0)$ and $(n,m)=(0,1)$ of O2 calculated using OE-WFAT and a bond length of $1.158$. The orbital used for this calculation is the HOMO, which is doubly degenerate and has a $\pi_{g}$ symmetry. The symmetry is reflected in the shape of the structure factor $|G_{00}|^{2}$ exhibiting two peaks at $\beta=42^{\circ}$ and $\beta=138^{\circ}$, and a mirror symmetry around $\beta=90^{\circ}$. The ionization yield of O2 interacting with an oscillating electric field of low intensity simulated using TDCI in Ref. [29] has a very similar double-peak structure located at $45^{\circ}$ and $135^{\circ}$. The shape as well as the position of these peaks also agrees with the experiments on O2 reported in Ref. [70]. The ME-WFAT results shown in Fig. 1(b) are very similar to the one-electron counterpart in panel (a) for both parabolic channels shown, with the only difference being the magnitude. In fact, the $r$ ratio for O2 obtained by HF is zero, with the largest coefficient belonging to HOMO, making the O2 Dyson orbital essentially the HOMO itself. The basis set, the ratio $r$, the largest and second-largest contributors to the Dyson orbital, and the ionization potentials used in ME-WFAT or OE-WFAT for O2 as well as other molecules studied here are summarized in Table 3.
Angle-dependent ionization yields of O2 have also been studied in the context of field-induced multielectron polarization effects (MEP) in Ref. [71]. It was found that MEP has negligible effects on the angle-dependent yield of O2. Although MEP is not considered in this work—remember that only field-free orbitals enter the formulation, this observation agrees with ours in that multielectron effects are negligible in the ionization from the ground state of O2 to that of its cation. To formally incorporate MEP in WFAT framework, one must include the 1st order correction of ME-WFAT, which, however, lies outside the scope of the present work, but will be studied in a future publication.
Table 3: The basis set, the $r$ ratio (Eq. (53)), the largest and 2nd largest neutral’s MO, and ionization potential (IP) of the molecules studied in this work. The IP’s are in hartree, and the values to the left and right of the slash sign is the IP for ME-WFAT and OE-WFAT, respectively.